OpenStudy (anonymous):

Given that n AM's are inserted between two sets of numbers a, 2b and 2a, b , where a and b are rational numbers. Suppose further that mth mean between these sets of number is same then the ratio of a:b equals

OpenStudy (anonymous):

@sathiprakash

OpenStudy (anonymous):

im sorry dude.....hope other helps you...im not in a mood now thanks

OpenStudy (anonymous):

@gerryliyana

OpenStudy (anonymous):

@mathslover

OpenStudy (anonymous):

@BostonBlue

OpenStudy (anonymous):

@Mashy

mathslover (mathslover):

See, I can give you a hint as I have to go for dinner

mathslover (mathslover):

Can you please wait for 15 minutes? I will help you then ...

OpenStudy (anonymous):

nevermind take your time

OpenStudy (anonymous):

@mathslover

OpenStudy (anonymous):

@BostonBlue got any idea

mathslover (mathslover):

try to go like this : 2b = a + nd b = 2a + nd' \(\cfrac{2b-a}{n} = d\) ------ (1) and \(\cfrac{b-2a}{n} = d'\) ----------(2) T_(m+1) = a + md... T' _ (m+1) = 2a + m/n (b-2a)

mathslover (mathslover):

@Meinme did you get what I meant?

mathslover (mathslover):

@Meinme are you there?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

i did it same but not getting the answer ILL check it once and come back